PART 2 Complete the following problems. 1. Because your 18 year-old friend never learned dimensional analysis, he started working at a fast food restaurant e works 5 days a week. He gets paid every 2 weeks with a salary of $440.34 (after taxes). pproximately how many hamburgers will he have to wrap to make his first one million dollars? Round to the nearest whole hamburger. Explicitly show all conversions in your setup. How much time will it take to wrap all those hamburgers (given as your answer to part a Express the time in years. Assume that he will work 40 weeks out of the year and that he wor with the same efficiency the nearest year b. every day. Also assume he works 40 hours/week. Round your answer to C. 2. Aluminum is a lightweight metal (density -2.70 g/cm) used in aircraft construction, high-voltage transmission lines, and foil. What is the density in kg/m3?
Answer #1

hamburgers per 3 hours= 350, no of workiing hours per day= 8 no of hamburgers per day = 350*8/3, no of days per two weeks = 2*5= 10 days, no of hamburgers for 10 days= 350*10*8/3

for two weeks( i.e 10 working days) he earns = $440.34

1 million= 10,00,000

440.34$ are for no of hamburgers of 350*10*8/3

10,00,000$ are to be earned by         10,00,000*350*10*8/(3*440.34)=21195743 hamburgers

350*10*8/3 hamburgers are wrapped in 2 weeks time

21195743 hamburgers are wrapped in 21195743*2/(350*10*8/3)=4542 weeks

there are 52 week in a years. Hence no of years = 4542/52=87 years

since he is 18 to-day, he would be 18+87= 105 years

2. density= 2.7 g/cc, 1 gm= 10-3 kg and 1cc= 10-6 m3

density= 2.7* 10-3/10-6 kg/m3=2700 kg/m3